OpenStudy (anonymous):

I am solving proportions, 6 over 1/3 = 1/2 over t, what is t? I tried it on my own and came up with t=1. Is that correct?

OpenStudy (anonymous):

6t=1/6 => t = 1/36

OpenStudy (anonymous):

can you please show me how you came to that answer? part of the insnstruction say to cross multiply to make sure that your answer is correct.

OpenStudy (anonymous):

6/(1/3) = 6 * 3

OpenStudy (anonymous):

6 is the numerator and 1/3 is the denominator = 1/2 is the numerator and "t" is the denominator. So i multiplied 1/3 and 1/2 = 1/6. To find "t" wouldnt 6/t = 1/6?

OpenStudy (anonymous):

6 is the numerator and 1/3 is the denominator = 1/2 is the numerator and "t" is the denominator means that 6/(1/3) = (1/2)/t => 6t = 1/2*1/3 = 1/6 => t= (1/6)/6 = 1/36

OpenStudy (anonymous):

use a calculator to be sure

OpenStudy (anonymous):

i came up with the same thing. Here is where I am confused. On every other problem I have done when cross multiplied they answers are the same number. Or =

OpenStudy (anonymous):

Nevermind, I got it. DUH! Thank you for your help.